# Measuring Spiritual Values and Bias of Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have become integral tool for users from various backgrounds. LLMs, trained on vast corpora, reflect the linguistic and cultural nuances embedded in their pre-training data. However, the values and perspectives inherent in this data can influence the behavior of LLMs, leading to potential biases. As a result, the use of LLMs in contexts involving spiritual or moral values necessitates careful consideration of these underlying biases. Our work starts with verification of our hypothesis by testing the spiritual values of popular LLMs. Experimental results show that LLMs’ spiritual values are quite diverse, as opposed to the stereotype of atheists or secularists. We then investigate how different spiritual values affect LLMs in social-fairness scenarios (e.g., hate speech identification). Our findings reveal that different spiritual values indeed lead to different sensitivity to different hate target groups. Furthermore, we propose to continue pre-training LLMs on spiritual texts, and empirical results demonstrate the effectiveness of this approach in mitigating spiritual bias.
@@
Warning: This paper may contain content some audiences find offensive or objectionable.
Measuring Spiritual Values and Bias of Large Language Models
Songyuan Liu1, Ziyang Zhang1, Runze Yan2, Wei Wu3, Carl Yang1, Jiaying Lu2† 1Department of Computer Science, Emory University, 2Center for Data Science, Nell Hodgson Woodruff School of Nursing, Emory University 3Department of Religion, Emory University Correspondence: Jiaying Lu, Ph.D., jiaying.lu@emory.edu
1 Introduction
@@
The popularity of Large Language Models (LLMs) Touvron et al. (2023); OpenAI (2024) has surged in recent years due to their remarkable capabilities in understanding natural language and assisting with humans to accomplish a wide range of tasks. These models are now deeply integrated into users’ daily workflow and entertainment, significantly influencing how individuals interact with digital systems in daily life. As LLMs continue to evolve, their widespread use has raised important questions about their inclusivity and neutrality, particularly when it comes to diverse user groups hold varied spiritual beliefs Riley-Smith and Throop (2022). It is essential to explore how LLMs handle religious and spiritual content. The inherent complexity of language generation Blodgett et al. (2020) in LLMs brings up concerns regarding potential biases, both subtle and overt, that could affect the spiritual values represented by these models.
@@
Throughout history, spiritual beliefs have emerged in nearly all human cultures Brown (2004); Murdock (1965); Johnson (2005). These beliefs evolved into religions Peoples et al. (2016), fostering group cooperation, human morals, and societal stability. However, differing beliefs in deities Smith (1991); Huntington (1996) often led to conflicts over religious supremacy Riley-Smith and Throop (2022). These rivalries have contributed to biases among individuals with different spiritual values in modern societies Esposito and Kalin (2011), manifesting as discrimination, hate speech, and cyber-abuse, which affect social dynamics.
@@
Humans hold different spiritual values because these beliefs are shaped by a variety of factors, including culture, tradition, upbringing, and personal experiences. Similarly, LLMs, while not conscious agents, may inherently reflect biases present in the vast amounts of data they are trained on Baeza-Yates (2018). Past studies have already highlighted the existence of ethical biases in LLMs Blodgett et al. (2020); Yoder et al. (2022), demonstrating how these models can reflect and amplify societal prejudices Feng et al. (2023). In this study, we first measure spiritual values of popular LLMs using two spirituality value assessments (Section 2). Experimental results show that LLMs’ spiritual values are very diverse, as opposed to our hypothesis that LLMs would tend to be more secular. We then quantify the effects of spiritual values on religion hate speech identification task (Section 3). Empirically, more religious LLMs tend to perform better in hate speech detection. This is also verified by our further pre-training experiment, where one language model achieves better performance after further unsupervised training on religion literature.
@@
Figure 1: Overview of the two spiritual value evaluation tools: SP-Typology and SP-10Axes.
2 Spiritual Values Evaluation
@@
Spiritual value assessment tools in the format of questionnaires are widely used in social science Pew Research Center (2019) and health science Saguil and Phelps (2012). Some pioneering work in the AI field also adopt domain specific questionnaires Feng et al. (2023) to assess certain social-oriented values of AI models. We follow a similar approach for our spiritual values evaluation.
2.1 Evaluation Setup
@@
Table 1: The statistics of the two assessments.
@@
Data Sources. Specifically, we adopt two questionnaire-style assessments tools “SP-Typology” test Pew Research Center (2019) and “SP-10Axes” bannnedb (2022) to measure the spiritual values of LLMs. Statistics of these two assessments are presented in Table 1. SP-Typology (Pew Center’s “religious typology”) categorizes individuals by shared spiritual beliefs, religious practice, and sources of meaning, unlike traditional assessments that group people by denomination. This system, independent of race, ethnicity, age, education, and political views, broadly divides people into three main groups, with finer distinctions in each: highly religious (Sunday Stalwarts, God-and-Country Believers, Diversely Devout), somewhat religious (Relaxed Religious, Spiritually Awake), and non-religious (Religion Resisters, Solidly Secular). SP-10Axes attempts to assign percentages on ten different religious value axes, by presenting 133 statements and collecting participant’s opinion (from strongly agree to strongly disagree) on these statements. After completing the assesment, participant will get the percentage on each of the ten axes: (1) Pro- / Anti-Catholic; (2) Pro- / Anti-Protestant; (3) Pro- / Anti-Orthodox; (4) Philo- / Anti-Semitic; (5) Islamophilic / Islamophobic; (6) Pro / Anti-Buddhist; (7) Pro / Anti-Hindu; (8) Pro / Anti-Pagan; (9) Satanic / Divine; (10) Atheistic / Religious. In our study, we treat LLMs as human participants to read the questions in the assessment and let them respond with their opinion by selecting their preferred answers. Toy examples of questions and prompts are shown in Figure 1.
@@
Evaluated LLMs. A diverse range of LLMs are evaluated: (a) “Commercial LLMs”: GPT-4-turbo OpenAI (2024). (b) “Open-source LLMs”: (RNN) RWKV-6 Peng et al. (2023); (State space model) Mamba-2.8b Gu and Dao (2024); (Transformer) LLAMA-2 Touvron et al. (2023), LLAMA-3 Llama Team (2024), Qwen-1.5-chat Qwen Team (2023), Phi-3 Microsoft (2024), Mistral Jiang et al. (2023), Gemma Team (2024), Vicuna-v1.5 Chiang et al. (2023), LongChat Li et al. (2023), FastChat-T5 Chiang et al. (2023), Tulu-2-dpo Ivison et al. (2023), Mpt-chat Team (2023), Chatglm2 GLM (2024). We use 7B size with default hyperparameters for most LLMs and greedy decoding for consistent generation.
Figure 2: SP-Typology Distribution Overview.
2.2 Evaluation Results and Analysis
@@
Figure 3: SP-10Axes Results of Selected LLMs. Lighter blue indicates value assessment towards more spiritual.
@@
We hypothesize that these LLMs would exhibit minimal expression of spiritual or religious values, largely due to the secular nature of the pre-training corpora, which predominantly consist of content from the Internet or publications known to be mostly devoid of religious or spiritual value preferences Shen et al. (2023); Byrd (2023); Elazar et al. (2024). Figure 2 present the overview of typology distribution on the SP-Typology, with a more detailed break-down in Appendix B. Contrary to our initial hypothesis that LLMs would exhibit secular tendencies, the findings reveal that the majority of the evaluated LLMs display either a somewhat or highly religious inclination. Notably, only one of the models are classified within the least religious category, "Solidly Secular." This observation suggests that the pre-training corpora of these LLMs contain a significant amount of spiritual or religious content, influencing the models to embody such values. On the other hand, the results of our experiment on the SP-10Axes are presented in Figure 3. For clarity and space considerations, we selected 10 representative LLMs for visualization. More experimental results can be found in Appendix C. The central vertical line in the figure denotes a neutral stance towards each category. Consistent with the observations from SP-Typology, the results reveal variability in the distribution of responses across the LLMs. For instance, models such as Vicuna and Tulu exhibit a predominantly neutral stance across all categories, while others, such as ChatGLM and Llama-2, display a more left-shifted (spiritually inclined) tendency. Furthermore, a closer inspection of individual LLMs reveals considerable variation in their consistency across different categories. While models like Vicuna and LongChat tend to maintain a neutral position around the 50 mark, others like Phi-3 and Llama-3 demonstrate more pronounced fluctuations in their scores across various categories. To specifically assess the spiritual inclinations of the LLMs, we focus on the final category, Atheistic vs. Religious. Our findings suggest that the LLMs generally lean towards a more religious or neutral stance. We further conduct a correlation analysis in Appendix 5, which demonstrates that individual model presents consistent spiritual values in both assessments.
@@
3 Quantifying the Spiritual Bias Effect on Social-Oriented Fairness Tasks
@@
3.1 Hate Speech Targeting Different Religions
We examines the influence on religion-targeted hate speech detection tasks. The hate-speech-identity dataset Yoder et al. (2022) (21,053 discourses) is adopted and we narrow the scope down into six major spiritual-oriented groups: “Judaism, Christianity, Islam, Hinduism, Buddhism, and Non-religious”. Details of the mapping between original identity groups and our spiritual-oriented groups are presented in Appendix E. Our evaluation uses F1 score to assess the LLMs’ performance in hate detection with regards to each spiritual group. For each evaluated instruction-tuned LLM, we configured the system prompt to position the LLM as an expert in detecting hate speech content. The output was binary, requiring the LLM to respond with either "Yes" or "No," for whether the input text contain hate speech towards a specific spiritual group. For our experiment, we choose one representative LLM from each religious group defined in Figure 4. The result is displayed in Table 2, which shows an inclination of performance in hate detection task if the LLM falls into a higher spiritual group.
Table 2: F1-score (in percentage) of selected LLMs on religion-hate speech detection.
@@ 3.1 Hate Speech Targeting Different Religions
Table 3: F1-score (in percentage) on religion-hate speech detection before and after further pre-training.
@@
GPT2 Pali 36.42 38.0 37.74 54.19 55.13 39.37 54.44
GPT2 Veda 41.68 45.41 43.09 66.60 69.50. 46.30 67.14
GPT2 Tanakh 36.34 46.17 36.91 59.95 60.23 44.0 59.86
3.2 Effects from Different training sources
@@ 4 Conclusion
While our study emphasizes the evaluation of spiritual values and biases in LLMs, we acknowledge several limitations in our approach. One key limitation is the nature of the assessment tools used, which may include behavior-based questions such as "How often do you attend religious services?" Since LLMs lack physical bodies and personal experiences, they can only respond to such questions in a hypothetical manner, which may limit the accuracy of their spiritual self-representation. Future work should consider developing assessment tools tailored specifically to the capabilities and limitations of LLMs in order to better capture their underlying biases in relation to spiritual and moral values. Additionally, LLMs’ responses may be shaped by the in-context learning examples and language used in the prompt, which could inadvertently influence their answers. auto-prompt generation techniques that minimize prompt bias and ensure more consistent and objective responses.
@@ Appendix A List of Evaluated LLMs
(a.1) GPT-4-turbo OpenAI (2024): a Transformer-based model designed by OpenAI, known for its conversational abilities and fine-tuning with RLHF.
@@
(b.1) RWKV-6 Peng et al. (2023): a unique model that utilizes a Recurrent Neural Network (RNN) architecture, offering efficiency while maintaining capabilities similar to Transformer-based models.
(b.2) LLAMA-2 Touvron et al. (2023): high-performing, Transformer-based models developed by Meta, notable for its efficient scaling and fine-tuning for a wide variety of tasks.
(b.3) LLAMA-3 Llama Team (2024): an improved version of LLAMA-2 by large traiing data, enhanced model architecture, improved hanlding of ambiguity and uncertainty, and enhanced safety features.
(b.4) Qwen-1.5-chat: a Transformer-based LLM developed by Alibaba, pretrained for up to 3 trillion tokens of multilingual data with a wide coverage of domains and languages.
(b.5) Phi-3 Microsoft (2024): a small yet powerful Transformer-based model developed by Microsoft.
(b.6) Mistral Jiang et al. (2023): a model that leverages grouped-query attention (GQA) for faster inference and sliding window attention (SWA) to handle sequences of arbitrary length.
(b.7) Gemma Team (2024): a light-weight model developed by Google AI.
(b.8) Vicuna-v1.5 Chiang et al. (2023): a model trained by fine-tuning llama-13b on user-shared conversations collected from ShareGPT.
@@
(b.11) Tulu-2-dpoIvison et al. (2023): a fine-tuned version of Llama 2 that was trained on a mix of publicly available, synthetic and human datasets.
@@
(b.13) Chatglm2 GLM (2024): a Transformer-based model fine-tuned for Chinese-language dialogue and language understanding tasks.
@@
Figure 4: Forest Map of SP-Typology Results.

### v2

# Measuring Spiritual Values and Biases of Large Language Models

@@ Abstract.
Large language models (LLMs) have become integral tool for users from various backgrounds. Pre-trained on vast corpora, LLMs reflect the linguistic and cultural nuances embedded in their training data. However, the values and perspectives inherent in this data can influence the behavior of LLMs, leading to potential biases. As a result, the use of LLMs in contexts involving spiritual or moral values necessitates careful consideration of these underlying biases. Our work starts with by testing the spiritual values of popular LLMs. Experimental results show that LLMs’ spiritual values are quite diverse, as opposed to the stereotype of atheists or secularists. We then investigate how different spiritual values affect LLMs in social-fairness scenarios (e.g., hate speech identification). Our findings reveal that different spiritual values indeed lead to varied sensitivity to different hate target groups. Furthermore, we propose to continue pre-training LLMs on spiritual texts, and empirical results demonstrate the effectiveness of this approach in mitigating spiritual biases.
@@
Warning: This paper contains contents some audiences may find offensive or objectionable.
conference: KDD’25 Workshop Large Language Models for Scientific and Societal Advances; August 3-7, 2025; Toronto, ON, Canada
1. Introduction
@@
The popularity of Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2024) has surged in recent years due to their remarkable capabilities in understanding natural language and assisting with humans to accomplish a wide range of tasks. These models are now deeply integrated into users’ daily workflow and entertainment, significantly influencing how individuals interact with digital systems in daily lives. As LLMs continue to evolve, their widespread use has raised important questions about their inclusivity and neutrality, particularly when it comes to diverse user groups holding varied spiritual beliefs (Riley-Smith and Throop, 2022). It is essential to explore how LLMs handle religious and spiritual content. The inherent complexity of language generation (Blodgett et al., 2020) in LLMs brings up concerns regarding potential biases, both subtle and overt, that could affect the spiritual values represented by these models.
@@
Throughout history, spiritual beliefs have emerged in nearly all human cultures (Brown, 2004; Murdock, 1965; Johnson, 2005). These beliefs evolved into religions (Peoples et al., 2016), fostering group cooperation, human morals, and societal stability. However, differing beliefs in deities (Smith, 1991; Huntington, 1996) often led to conflicts over religious supremacy (Riley-Smith and Throop, 2022). These rivalries have contributed to biases among individuals with different spiritual values in modern societies (Esposito and Kalin, 2011), manifesting as discrimination, hate speech, and cyber-abuse, which affect social dynamics.
@@
Humans hold different spiritual values because these beliefs are shaped by a variety of factors, including culture, tradition, upbringing, and personal experiences. Similarly, LLMs, although as unconscious agents, may inherently reflect biases present in the vast amounts of data they are trained on (Baeza-Yates, 2018). Past studies have already highlighted the existence of ethical biases in LLMs (Blodgett et al., 2020; Yoder et al., 2022), demonstrating how these models can reflect and amplify societal prejudices (Feng et al., 2023). In this study, we first measure spiritual values of popular LLMs using two spirituality value assessments (Section 2). Experimental results show that LLMs’ spiritual values are very diverse, as opposed to our hypothesis that LLMs would tend to be more secular. We then quantify the effects of spiritual values on religion hate speech identification task (Section 3). Empirically, more religious LLMs tend to perform better in hate speech detection. This is also verified by our further pre-training experiments, where one language model achieves better performance after further unsupervised training on religion literature.
@@
Figure 1. Overview of the two spiritual value evaluation tools: SP-Typology and SP-10Axes.
2. Spiritual Values Evaluation
@@
Spiritual value assessment tools in the format of questionnaires are widely used in social science (Pew Research Center, 2019) and health science (Saguil and Phelps, 2012). Some pioneering work in the AI field also adopt domain specific questionnaires (Feng et al., 2023) to assess certain social-oriented values of AI models. We follow a similar approach for our spiritual values evaluation.
2.1. Evaluation Setup
@@
Table 1. The statistics of the two assessments.
@@
Data Sources. Specifically, we adopt two questionnaire-style assessments tools “SP-Typology” test (Pew Research Center, 2019) and “SP-10Axes” (bannnedb, 2022) to measure the spiritual values of LLMs. Statistics of these two assessments are presented in Table 1. SP-Typology (Pew Center’s “religious typology”) categorizes individuals by shared spiritual beliefs, religious practice, and sources of meaning, unlike traditional assessments that group people by denomination. This system, independent of race, ethnicity, age, education, and political views, broadly divides people into three main groups, with finer distinctions in each: highly religious (Sunday Stalwarts, God-and-Country Believers, Diversely Devout), somewhat religious (Relaxed Religious, Spiritually Awake), and non-religious (Religion Resisters, Solidly Secular). SP-10Axes attempts to assign percentages on ten different religious value axes, by presenting 133 statements and collecting participant’s opinion (from strongly agree to strongly disagree) on these statements. After completing the assesment, participant will get the percentage on each of the ten axes: (1) Pro- / Anti-Catholic; (2) Pro- / Anti-Protestant; (3) Pro- / Anti-Orthodox; (4) Philo- / Anti-Semitic; (5) Islamophilic / Islamophobic; (6) Pro / Anti-Buddhist; (7) Pro / Anti-Hindu; (8) Pro / Anti-Pagan; (9) Satanic / Divine; (10) Atheistic / Religious. In our study, we treat LLMs as human participants to read the questions in the assessment and let them respond with their opinion by selecting their preferred answers. Toy examples are shown in Figure 1.
@@
Evaluated LLMs. A diverse range of LLMs are evaluated: (a) “Commercial LLMs”: GPT-4-turbo (OpenAI, 2024). (b) “Open-source LLMs”: (RNN) RWKV-6 (Peng et al., 2023); (State space model) Mamba-2.8b (Gu and Dao, 2024); (Transformer) LLAMA-2 (Touvron et al., 2023), LLAMA-3 (Llama Team, 2024), Qwen-1.5-chat (Qwen Team, 2023), Phi-3 (Microsoft, 2024), Mistral (Jiang et al., 2023), Gemma (Team, 2024), Vicuna-v1.5 (Chiang et al., 2023), LongChat (Li et al., 2023), FastChat-T5 (Chiang et al., 2023), Tulu-2-dpo (Ivison et al., 2023), Mpt-chat (Team, 2023), Chatglm2 (GLM, 2024). We use 7B size with default hyperparameters for most LLMs and greedy decoding for consistent generation.
Figure 2. SP-Typology Distribution Overview.
2.2. Evaluation Results and Analysis
@@
Figure 3. SP-10Axes Results of Selected LLMs. Lighter blue indicates value assessment towards more spiritual.
@@
We hypothesize that these LLMs would exhibit minimal expression of spiritual or religious values, largely due to the secular nature of the pre-training corpora, which predominantly consist of content from the Internet or publications known to be mostly devoid of religious or spiritual value preferences (Shen et al., 2023; Byrd, 2023; Elazar et al., 2024). Figure 2 present the overview of typology distribution on the SP-Typology, with a more detailed break-down in Appendix B. Contrary to our initial hypothesis that LLMs would exhibit secular tendencies, the findings reveal that the majority of the evaluated LLMs display either a somewhat or highly religious inclination. Notably, only one of the models are classified within the least religious category, "Solidly Secular." This observation suggests that the pre-training corpora of these LLMs contain a significant amount of spiritual or religious content, influencing the models to embody such values. On the other hand, the results of our experiment on the SP-10Axes are presented in Figure 3. For clarity and space considerations, we selected 10 representative LLMs for visualization. More experimental results can be found in Appendix C. The central vertical line in the figure denotes a neutral stance towards each category. Consistent with the observations from SP-Typology, the results reveal variability in the distribution of responses across the LLMs. For instance, models such as Vicuna and Tulu exhibit a predominantly neutral stance across all categories, while others, such as ChatGLM and Llama-2, display a more left-shifted (spiritually inclined) tendency. Furthermore, a closer inspection of individual LLMs reveals considerable variation in their consistency across different categories. While models like Vicuna and LongChat tend to maintain a neutral position around the 50 mark, others like Phi-3 and Llama-3 demonstrate more pronounced fluctuations in their scores across various categories. To specifically assess the spiritual inclinations of the LLMs, we focus on the final category, Atheistic vs. Religious. Our findings suggest that the LLMs are generally inclined to a religious or neutral stance.
Figure 4. Correlation between the SP-Typology and SP-10Axes
Although we rigorously assess the spiritual values of LLMs using similar formats, due to the different institutional origins of these assessments, their objectives are inherently distinct. Specifically, the SP-Typology aims to measure the degree of religiosity in assessments’ takers, while the SP-10Axes seeks to categorize individuals into various religious groups. Consequently, it is challenging to assert that LLMs exhibit aligned spiritual values based solely on a direct comparison of the assessment results.
@@
3. Quantifying the Spiritual Bias Effect on Social-Oriented Fairness Tasks
@@
3.1. Hate Speech Targeting Different Religions
We examines the influence on religion-targeted hate speech detection tasks. The hate-speech-identity dataset (Yoder et al., 2022) (21,053 discourses) is adopted and we narrow the scope down into six major spiritual-oriented groups: “Judaism, Christianity, Islam, Hinduism, Buddhism, and Non-religious”. Details of the mapping method from the dataset’s original identity groups to our spiritual-oriented groups are presented in Appendix D. Our evaluation uses F1 score to assess the LLMs’ performance in hate detection with regards to each spiritual group. For each evaluated instruction-tuned LLM, we configured the system prompt to position the LLM as an expert in detecting hate speech content. The output was binary, requiring the LLM to respond with either "Yes" or "No," for whether the input text contain hate speech towards a specific spiritual group. For our experiment, we choose one representative LLM from each religious group defined in Figure 5. The result is displayed in Table 2, which shows an inclination of performance in hate detection task if the LLM falls into a higher spiritual group.
Table 2. F1-score (in percentage) of selected LLMs on religion-hate speech detection.
@@ 3.1. Hate Speech Targeting Different Religions
Table 3. F1-score (in percentage) on religion-hate speech detection before and after further pre-training.
@@
GPT2 Pali 36.42 38.00 37.74 54.19 55.13 39.37 54.44
GPT2 Veda 41.68 45.41 43.09 66.60 69.50 46.30 67.14
GPT2 Tanakh 36.34 46.17 36.91 59.95 60.23 44.00 59.86
3.2. Effects from Different training sources
@@ 4. Conclusion
While our study emphasizes the evaluation of spiritual values and biases in LLMs, we acknowledge several limitations in our approach. One key limitation is the nature of the assessment tools used, which may include behavior-based questions such as "How often do you attend religious services?" Since LLMs lack physical bodies and personal experiences, they can only respond to such questions in a hypothetical manner, which may limit the accuracy of their spiritual self-representation. Future work should consider developing assessment tools tailored specifically to the capabilities and limitations of LLMs in order to better capture their underlying biases in relation to spiritual and moral values. Additionally, LLMs’ responses may be shaped by the in-context learning examples and language used in the prompt, which could inadvertently influence their answers. In the future, auto-prompt generation techniques that minimize prompt bias can be used to ensure more consistent and objective responses.
@@
Acknowledgment
Research reported in this work was partially supported by the Atlanta Interdisciplinary Artificial Intelligence Network’s 2024 Seed Grant. The content is solely the responsibility of the authors and does not necessarily represent the official views of the the Atlanta Interdisciplinary Artificial Intelligence Network.
@@ Appendix A List of Evaluated LLMs
(a.1) GPT-4-turbo (OpenAI, 2024): a Transformer-based model designed by OpenAI, known for its conversational abilities and fine-tuning with RLHF.
@@
(b.1) RWKV-6 (Peng et al., 2023): a unique model that utilizes a Recurrent Neural Network (RNN) architecture, offering efficiency while maintaining capabilities similar to Transformer-based models.
(b.2) LLAMA-2 (Touvron et al., 2023): high-performing, Transformer-based models developed by Meta, notable for its efficient scaling and fine-tuning for a wide variety of tasks.
(b.3) LLAMA-3 (Llama Team, 2024): an improved version of LLAMA-2 by large traiing data, enhanced model architecture, improved hanlding of ambiguity and uncertainty, and enhanced safety features.
(b.4) Qwen-1.5-chat (Qwen Team, 2023): a Transformer-based LLM developed by Alibaba, pretrained for up to 3 trillion tokens of multilingual data with a wide coverage of domains and languages.
(b.5) Phi-3 (Microsoft, 2024): a small yet powerful Transformer-based model developed by Microsoft.
(b.6) Mistral (Jiang et al., 2023): a model that leverages grouped-query attention (GQA) for faster inference and sliding window attention (SWA) to handle sequences of arbitrary length.
(b.7) Gemma (Team, 2024): a light-weight model developed by Google AI.
(b.8) Vicuna-v1.5 (Chiang et al., 2023): a model trained by fine-tuning llama-13b on user-shared conversations collected from ShareGPT.
@@
(b.11) Tulu-2-dpo(Ivison et al., 2023): a fine-tuned version of Llama 2 that was trained on a mix of publicly available, synthetic and human datasets.
@@
(b.13) Chatglm2 (GLM, 2024): a Transformer-based model fine-tuned for Chinese-language dialogue and language understanding tasks.
@@
Figure 5. Forest Map of SP-Typology Results.